OpenStudy (math_bender_15):

A line passes through the point (-6, -8) and has a slope of 4 . Write an equation for this line.

OpenStudy (math_bender_15):

@tkhunny

OpenStudy (tkhunny):

Do you have a point-slope form?

OpenStudy (math_bender_15):

do you mean like y=mx+b?

OpenStudy (tkhunny):

No, that is the slope-intercept form. You should also have a point-slope form. Do you?

OpenStudy (tkhunny):

Usually, \((y-y_{0}) = m(x-x_{0})\)